# Rational Design Heterobilayers Photocatalysts for Efficient Water Splitting Based on 2D Transition-Metal Dichalcogenide and Their Janus

### v1

@@ 3.3 Solar to Hydrogen Conversion Efficiency (STH)
After screening all possible materials for their ability to initialize water splitting, we identified 9 materials (see Table 2) with an effective band gap greater than 1.23 eV, indicating that these materials can potentially straddle the whole redox potential of H2O. The redox potential of water depends on pH and can be described as follows:
$E_{\text{H}^{+}/\text{H}_{2}}=0-0.059\times\text{pH},$ (2)
@@
$E_{\text{O}_{2}/\text{H}_{2}\text{O}}=1.23-0.059\times\text{pH}.$ (3)
Determining the appropriate pH for water splitting is crucial for optimizing the performance of photocatalysts. Therefore, we analyzed the band alignment of the selected materials for the redox potentials at $\text{pH}=0$ and $7$, as illustrated in Figure S3 (Supporting Information). We find that some materials require solution $\text{pH}<-2$, indicating highly acidic conditions, which limits their practical application 52, 53. Therefore, these materials are excluded from the STH efficiency calculations. Only materials capable of splitting water at $\text{pH}=0$ and $7$ are considered for the STH efficiency evaluation, with the relevant material information provided in Table 2. It is noted that the VBM of WSe2-SWSe is slightly higher than the OER potential. Therefore, at $\text{pH}=7$, this material exhibits only the hydrogen evolution half-reaction during photocatalysis.
@@
WSe2-SWSe 0.1 0.01 96.63% 45.50% 30.71%
WSe2-TeWSe 0.12 0.25 87.08% 46.45% 30.72%
WS2-SMoSe 0.09 0.39 88.54% 52.77% 33.24%
WS2-SeMoS 0.27 0.73 60.68% 57.62% 29.12%
@@
The STH efficiency $\eta_{\text{STH}}$ of the selected materials were calculated using the following expression 1:
@@
$\eta_{\mathrm{STH}}=\eta_{\mathrm{abs}}\times\eta_{\mathrm{cu}}\times\frac{\left.\int_{0}^{\infty}P(h\omega)\mathrm{d}(h\omega)\right)}{\int_{0}^{\infty}P(h\omega)\mathrm{d}(h\omega)+\Delta\Phi\int_{E_{\mathrm{g}}}^{\infty}\frac{P(h\omega)}{h\omega}\mathrm{d}(h\omega)},$ (4)
@@
where $\eta_{\text{abs}}$ is the light absorption efficiency, $\eta_{\text{cu}}$ is the carrier utilization efficiency, $E_{g}$ is the band gap energy, $\Delta\Phi$ is the electric potential difference, and $P(h\omega)$ represents the AM1.5G solar flux at photon energy $h\omega$. The light absorption efficiency ($\eta_{\text{abs}}$) is given by
$\eta_{\mathrm{abs}}=\frac{\int_{E_{\mathrm{g}}}^{\infty}P(h\omega)\mathrm{d}(h\omega)}{\int_{0}^{\infty}P(h\omega)\mathrm{d}(h\omega)},$ (5)
and the carrier utilization $\eta_{\text{cu}}$ is determined by
@@
$\eta_{\mathrm{cu}}=\frac{\Delta E_{\text{spliting}}\int_{E_{\min}}^{\infty}\frac{P(h\omega)}{h\omega}\mathrm{d}(h\omega)}{\int_{E_{\mathrm{g}}}^{\infty}P(h\omega)\mathrm{d}(h\omega)}$ (6)
where $\Delta E_{\text{spliting}}$ = 1.23 eV for water splitting and
$E_{\text{min }}=\left\{\begin{array}[]{cc}\mathrm{E}_{\mathrm{g}}&\text{ for }\chi\left(\mathrm{H}_{2}\right)\geq 0.2,\chi\left(\mathrm{O}_{2}\right)\geq 0.6\\ E_{\mathrm{g}}+0.2-\chi\left(\mathrm{H}_{2}\right)&\text{ for }\chi\left(\mathrm{H}_{2}\right)<0.2,\chi\left(\mathrm{O}_{2}\right)\geq 0.6\\ E_{\mathrm{g}}+0.6-\chi\left(\mathrm{O}_{2}\right)&\text{ for }\chi\left(\mathrm{H}_{2}\right)\geq 0.2,\chi\left(\mathrm{O}_{2}\right)<0.6\\ E_{\mathrm{g}}+0.8-\chi\left(\mathrm{H}_{2}\right)-\chi\left(\mathrm{O}_{2}\right)&\text{ for }\chi\left(\mathrm{H}_{2}\right)<0.2,\chi\left(\mathrm{O}_{2}\right)<0.6\end{array}\right.,$ (7)
where $\chi(\text{H}_{2})$ and $\chi(\text{O}_{2})$ are the overpotentials for the hydrogen and oxygen evolution reactions, respectively. It is noted that $\chi(\text{H}_{2})$ is the potential difference between the CBM and the redox potential of $\text{H}^{+}/\text{H}_{2}$, while $\chi(\text{O}_{2})$ is the potential difference between the VBM and the redox potential of $_{2}/\text{H}_{2}\text{O}$ O, as shown in Figure 2. Our calculated results, presented in Table 2, show that the selected materials exhibit remarkable photocatalytic efficiencies and exceed the conventional theoretical limit of 18% 54. In particular, WS2-SMoSe achieves an $\eta_{\text{STH}}$ of 33.24%. In comparison, materials reported in previous studies, such as Penta-PdSeTe (29.71%) 55, Janus $\gamma$-GeSSe (28.78%) 1, and SPtGeAs2 (30.74%) 56, show competitive but lower efficiencies.
@@
The exceptional performance of WS2-SMoSe and other materials in Table 2 can be attributed to its optimal electronic structure and favorable band alignment, which facilitate efficient light absorption and charge carrier utilization. The light absorption efficiency $\eta_{\text{abs}}$ is particularly high for materials with smaller band gaps, as shown in Equation 5. WSe2-SWSe with a band gap of 0.604 eV exhibits an outstanding $\eta_{\text{abs}}$ of 96.63%, which allows it to absorb a broader spectrum of solar radiation compared to traditional photocatalysts with a standard band gap of 1.23 eV. It is important to note that Equation 5 assumes the absorbance as a step function zero for photon energies below the band gap and unity for energies above it. This simplification may lead to underestimating $\eta_{\text{abs}}$, particularly at lower photon energies. Therefore, while the calculated $\eta_{\text{abs}}$ provides valuable insights, it may be more accurate for materials with high photon energy absorption.
3.4 Carrier mobility
@@
The relationship between carrier mobility and photocatalytic efficiency is essential in determining the performance of photocatalysts. After the electron-hole pairs are separated, the free photoexcited carriers need to migrate to the chemical active sites for water splitting. Higher mobility enables these charge carriers to travel faster and further, reducing the probability of recombination before they can reach the reactive surface sites 57, 58, 59, 60. Mobility is primarily limited by phonon scattering, which impedes the movement of charge carriers. Traditional approaches to calculating mobility, such as deformation potential theory (DPT), focus on the scattering caused by longitudinal acoustic (LA) phonons 14, 60, 56. However, recent studies, such as Cheng and Liu have demonstrated that this method often underestimates the influence of longitudinal optical (LO) phonon scattering, particularly in Janus materials with strong polar effects.
@@
The carrier mobility $\mu_{\text{LA}}$ due to the LA phonon scattering within the DPT theory is given as follows 59, 1:
@@
where the effective masses along the zigzag ($x$) and armchair ($y$) directions are represented by $m_{x}^{*}$ and $m_{y}^{*}$, respectively, and the average effective mass given by $m_{d}=\sqrt{m_{x}^{*}m_{y}^{*}}$. A $\sqrt{3}\times 1$ supercell (see Figure S4 in Supporting Information) is used to compute $m_{x}^{*}$ and $m_{y}^{*}$. The in-plane Young’s modulus, $C_{2\mathrm{D}}$, is defined as 59, 1:
@@
Table 3: Material properties of selected Janus-TMDC heterobilayers, including effective masses ($m^{*}$), deformation potentials ($E_{d})$, and calculated carrier mobilities ($\mu_{\text{LA}}$).
Sample Direction Carrier $m*/m_{0}$ $E_{d}$ $C_{\text{2D}}$ $\mu_{\text{LA}}$
[…]
WSe2-SWSe zigzag h 0.401 2.26 218.7755 5085.70
e 0.225 2.70 11120.74
armchair h 0.499 2.18 201.1680 4038.86
e 0.290 2.64 8298.47
WSe2-TeWSe zigzag h 0.611 2.52 212.9836 1735.10
e 0.257 2.16 13274.89
armchair h 0.743 2.46 199.0206 1399.13
e 0.316 2.26 9215.51
WS2-SMoSe zigzag h 0.252 4.70 245.0984 3499.11
e 0.343 4.70 1697.92
armchair h 0.285 4.78 229.1263 2796.32
e 0.480 4.34 1330.22
[…]
Table 4: Calculated mobilities contributed by LO1 and LO2 modes ($\mu_{\text{LO1}}$ and $\mu_{\text{LO2}}$) and combined LA and LO modes ($\mu_{\text{total}}$).
Sample Direction Carrier $\mu_{\text{LO1}}$ $\mu_{\text{LO2}}$ $\mu_{\text{total}}$
[…]
WSe2-SWSe zigzag h 980.78 5669.29 718.07
e 1719.25 9937.93 1295.01
armchair h 980.78 5669.29 692.72
e 1719.25 9937.93 1245.67
WSe2-TeWSe zigzag h 946.32 246.97 175.99
e 2237.96 584.06 447.56
armchair h 946.32 246.97 171.81
e 2237.96 584.06 441.01
WS2-SMoSe zigzag h 109174.17 2137.68 1311.06
e 72065.71 1411.08 762.48
armchair h 109174.17 2137.68 1198.22
e 72065.71 1411.08 678.28
It is important to note that many 2D TMDCs exhibit mobility limited by not only LA but also LO phonon scattering 28. The LO phonon scattering is closely linked to the Born effective charge, which reflects the degree of electric polarization change due to atomic vibrations. Thus, a large Born effective charge correlates with stronger carrier scattering, leading to reduced mobility. This critical insight highlights the need to go beyond DPT to accurately assess the intrinsic mobility of Janus-TMDC heterostructures when LO phonon scattering could play a substantial role. In the present study, the Fröhlich interaction model 28, 30 is employed to quantify the mobility contributions from the LO phonon mode. The total scattering rate near the band edges can be written as
@@
$\begin{split}\frac{1}{\tau_{\text{LO}}}&=\frac{e^{2}n_{h\omega}}{M_{M}M_{X1}M_{X2}S_{0}d_{eff}^{2}(\varepsilon\varepsilon_{0})^{2}(\hbar\omega)^{2}}\times\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left[\left|Z_{MB}\right|\sqrt{M_{X1}M_{X2}}+\left|Z_{X1}\right|\sqrt{M_{M}M_{X2}}+\left|Z_{X2}\right|\sqrt{M_{M}M_{X1}}\right]^{2},\end{split}$ (11)
@@
where $e$ is the elementary charge, $n_{h\omega}$ is the Bose-Einstein distribution at the phonon frequency at the gamma point, $A$ is the equilibrium area, $d_{eff}$ is the effective thickness, $\varepsilon$ is the in-plane optical dielectric constant, and $\varepsilon_{0}$ is the dielectric constant. Additionally, $M_{M}$, $M_{X1}$, and $M_{X2}$ are the atomic masses of the metal and the two chalcogens in the structure, while $Z_{MB}$, $Z_{X1}$, and $Z_{X2}$ are the Born effective charges of the metal and the two chalcogen atoms (see Table S2 in Supporting Information for reference). Thus, the mobility contributed by the LO phonon mode is given by
@@
In Janus-TMDC heterobilayers, two distinct LO phonon modes, LO1 and LO2, correspond to the vibrations of the TMDC and Janus layers, respectively, as shown in Figure 3. The calculated mobilities for $\mu_{\text{LO1}}$ and $\mu_{\text{LO2}}$ are provided in Table 4. Then the total mobility $\mu_{\text{total}}$ is then determined using Matthiessen’s rule as follows 28:
@@
The calculated $\mu_{\text{LO1}}$, $\mu_{\text{LO2}}$, and $\mu_{\text{total}}$ are listed in Table  4. Our results reveal that LO phonon scattering significantly limits mobility in Janus-TMDC heterobilayers, which leads to lower carrier mobility compared to those predicted by DPT alone (see Table  3).

### v2

@@ 3.4 Solar to Hydrogen Conversion Efficiency (STH)
After screening all possible materials for their ability to initialize water splitting, we identified 9 materials (see Table 2) with an effective band gap greater than 1.23 eV, indicating that these materials can potentially straddle the whole redox potential of H2O. The redox potential of water depends on pH and can be described as follows 43, 44:
$E_{\text{H}^{+}/\text{H}_{2}}=-4.44+0.059\times\text{pH},$ (4)
@@
$E_{\text{O}_{2}/\text{H}_{2}\text{O}}=-5.67+0.059\times\text{pH}.$ (5)
Determining the appropriate pH for water splitting is crucial for optimizing the performance of photocatalysts. Therefore, we analyzed the band alignment of the selected materials for the redox potentials at $\text{pH}=0$ and $7$, as illustrated in Figure S4 (Supporting Information). We find that some materials require solution $\text{pH}<-2$, indicating highly acidic conditions, which limits their practical application 45, 46. Therefore, these materials are excluded from the STH efficiency calculations. Only materials capable of splitting water at $\text{pH}=0$ and $7$ are considered for the STH efficiency evaluation, with the relevant material information provided in Table 2. It is noted that the VBM of WSe2-SWSe is slightly higher than the OER potential. Therefore, at $\text{pH}=7$, this material exhibits only the hydrogen evolution half-reaction during photocatalysis. Interestingly, all the materials identified as suitable for photocatalytic water splitting exhibit a Z-scheme alignment. From our perspective, this suggests that a large percentage of viable photocatalysts in this group of Janus/TMDC heterobilayers adopt the Z-scheme configuration. This can be attributed to their effective band alignment, which enables direct electron transfer between semiconductors with complementary band structures, making them highly suitable for photocatalytic applications.
@@
WSe2-SWSe 0.10 0.01 96.63% 22.75% 15.35%
WSe2-TeWSe 0.12 0.25 87.08% 23.23% 15.36%
WS2-SMoSe 0.09 0.39 88.54% 26.39% 16.62%
WS2-SeMoS 0.27 0.73 60.68% 28.81% 14.56%
@@
Table 3: Material properties of selected Janus-TMDC heterobilayers, including effective masses ($m^{*}$), deformation potentials ($E_{d})$, calculated carrier mobilities ($\mu_{\text{LA}}$), calculated mobilities contributed by LO1 and LO2 modes ($\mu_{\text{LO1}}$ and $\mu_{\text{LO2}}$) and combined LA and LO modes ($\mu_{\text{total}}$).
Sample Direction Carrier $m*/m_{0}$ $E_{d}$ $C_{\text{2D}}$ $\mu_{\text{LA}}$ $\mu_{\text{LO1}}$ $\mu_{\text{LO2}}$ $\mu_{\text{total}}$
[…]
WSe2-SWSe zigzag h 0.401 2.26 218.78 5085.70 980.78 5669.29 718.07
e 0.225 2.70 11120.74 1719.25 9937.93 1295.01
armchair h 0.499 2.18 201.17 4038.86 980.78 5669.29 692.72
e 0.290 2.64 8298.47 1719.25 9937.93 1245.67
WSe2-TeWSe zigzag h 0.611 2.52 212.98 1735.10 946.32 246.97 175.99
e 0.257 2.16 13274.89 2237.96 584.06 447.56
armchair h 0.743 2.46 199.02 1399.13 946.32 246.97 171.81
e 0.316 2.26 9215.51 2237.96 584.06 441.01
WS2-SMoSe zigzag h 0.252 4.70 245.10 3499.11 109174.17 2137.68 1311.06
e 0.343 4.70 1697.92 72065.71 1411.08 762.48
armchair h 0.285 4.78 229.13 2796.32 109174.17 2137.68 1198.22
e 0.480 4.34 1330.22 72065.71 1411.08 678.28
@@
The STH efficiency $\eta_{\text{STH}}$ of the selected materials were calculated using the following expression 1, 47:
@@
$\eta_{\mathrm{STH}}=\eta_{\mathrm{abs}}\times\eta_{\mathrm{cu}}\times\frac{\left.\int_{0}^{\infty}P(\hbar\omega)\mathrm{d}(\hbar\omega)\right)}{\int_{0}^{\infty}P(\hbar\omega)\mathrm{d}(\hbar\omega)+\Delta\Phi\int_{E_{\mathrm{g}}}^{\infty}\frac{P(\hbar\omega)}{\hbar\omega}\mathrm{d}(\hbar\omega)},$ (6)
@@
where $\eta_{\text{abs}}$ is the light absorption efficiency, $\eta_{\text{cu}}$ is the carrier utilization efficiency, $E_{g}$ is the band gap energy, $\Delta\Phi$ is the electric potential difference, and $P(\hbar\omega)$ represents the AM1.5G solar flux at photon energy $\hbar\omega$. The light absorption efficiency ($\eta_{\text{abs}}$) is given by:
$\eta_{\mathrm{abs}}=\frac{\int_{E_{\mathrm{g}}}^{\infty}P(\hbar\omega)\mathrm{d}(\hbar\omega)}{\int_{0}^{\infty}P(\hbar\omega)\mathrm{d}(\hbar\omega)},$ (7)
, the carrier utilization $\eta_{\text{cu}}$ is determined by:
@@
$\eta_{\mathrm{cu}}=\frac{1}{2}\times\frac{\Delta E_{\text{spliting}}\int_{E_{\min}}^{\infty}\frac{P(\hbar\omega)}{\hbar\omega}\mathrm{d}(\hbar\omega)}{\int_{E_{\mathrm{g}}}^{\infty}P(\hbar\omega)\mathrm{d}(\hbar\omega)},$ (8)
where $\Delta E_{\text{spliting}}$ = 1.23 eV for water splitting and $E_{\text{min}}$ is defined as
$E_{\text{min}}=\begin{cases}0&\text{for}~{}~{}\chi\left(\mathrm{H}_{2}\right)\geq 0.2,\chi\left(\mathrm{O}_{2}\right)\geq 0.6\\ E_{\mathrm{g}}+0.2-\chi\left(\mathrm{H}_{2}\right)&\text{for}~{}~{}\chi\left(\mathrm{H}_{2}\right)<0.2,\chi\left(\mathrm{O}_{2}\right)\geq 0.6\\ E_{\mathrm{g}}+0.6-\chi\left(\mathrm{O}_{2}\right)&\text{for}~{}~{}\chi\left(\mathrm{H}_{2}\right)\geq 0.2,\chi\left(\mathrm{O}_{2}\right)<0.6\\ E_{\mathrm{g}}+0.8-\chi\left(\mathrm{H}_{2}\right)-\chi\left(\mathrm{O}_{2}\right)&\text{for}~{}~{}\chi\left(\mathrm{H}_{2}\right)<0.2,\chi\left(\mathrm{O}_{2}\right)<0.6\end{cases},$ (9)
where $\chi(\text{H}_{2})$ and $\chi(\text{O}_{2})$ are the overpotentials for the hydrogen and oxygen evolution reactions, respectively. It is noted that $\chi(\text{H}_{2})$ is the potential difference between the CBM and the redox potential of $\text{H}^{+}/\text{H}_{2}$, while $\chi(\text{O}_{2})$ is the potential difference between the VBM and the redox potential of $_{2}/\text{H}_{2}\text{O}$ O, as shown in Figure 2. The factor 1/2 in the carrier utilization efficiency formula implies that only half of the photoinduced carriers (electrons and holes) are effectively utilized for the HER and OER in the Z-scheme system. This is because the Z-scheme mechanism involves the recombination of the electrons from the CBM of one material with the holes in the VBM of another material at the interface 47. In particular, WS2-SMoSe achieves the highest $\eta_{\text{STH}}$ of 16.62%. This demonstrates a significant improvement compared to previously reported Z-scheme vdW materials. For example, MoSe2/Ti2CO exhibits an STH efficiency of 12% 48, while MoSe2/SnSe2 and WSe2/SnSe2 achieve efficiencies of 10.5% and 9.3%, respectively 49. Additionally, CrS3/GeSe achieves 10.60% 50, and ZrS2/SeMoS reaches 6.3% 51. Furthermore, the predicted STH efficiency surpasses the critical threshold of 10% required for commercial applications, highlighting its potential as a promising photocatalyst for hydrogen production.
@@
Although the efficiency of the Z-scheme is only half that of Type II or Type I alignments, its importance becomes evident when the intrinsic band gap of the materials narrows. The Z-scheme enables these materials to function as photocatalysts by aligning their VBM and CBM to straddle both the HER and OER potentials, as reported in Wang et al.’s works. The promising performance of WS2-SMoSe and other materials in Table 2 can be attributed to its optimal electronic structure and favorable band alignment, which facilitate efficient light absorption and charge carrier utilization. The light absorption efficiency $\eta_{\text{abs}}$ is particularly high for materials with smaller band gaps, as shown in Eq. 7. WSe2-SWSe with a band gap of 0.604 eV exhibits an outstanding $\eta_{\text{abs}}$ of 96.63%, which allows it to absorb a broader spectrum of solar radiation compared to traditional photocatalysts with a standard band gap of 1.23 eV.
It is important to note that Eq. 7 assumes the absorbance as a step function zero for photon energies below the band gap and unity for energies above it. This simplification may lead to underestimating $\eta_{\text{abs}}$, particularly at lower photon energies. Therefore, while the calculated $\eta_{\text{abs}}$ provides valuable insights, it may be more accurate for materials with high photon energy absorption.
[…]
3.5 Precision and efficiency approach in transport calculation
@@
The relationship between carrier mobility and photocatalytic efficiency is essential in determining the performance of photocatalysts. After the electron-hole pairs are separated, the free photoexcited carriers need to migrate to the chemical active sites for water splitting. Higher mobility enables these charge carriers to travel faster and further, reducing the probability of recombination before they can reach the reactive surface sites 52, 53, 54, 55. Mobility is primarily limited by phonon scattering, which impedes the movement of charge carriers. Traditional approaches to calculating mobility, such as deformation potential theory (DPT), focus on the scattering caused by longitudinal acoustic (LA) phonons 8, 55, 1. However, Cheng and Liu 25 have demonstrated that this method often underestimates the influence of longitudinal optical (LO) phonon scattering, particularly in Janus materials with strong polar effects.
@@
The carrier mobility $\mu_{\text{LA}}$ due to the LA phonon scattering within the DPT theory is given as follows 55, 1:
@@
where the effective masses along the zigzag ($x$) and armchair ($y$) directions are represented by $m_{x}^{*}$ and $m_{y}^{*}$, respectively, and the average effective mass given by $m_{d}=\sqrt{m_{x}^{*}m_{y}^{*}}$. A $\sqrt{3}\times 1$ supercell (see Figure S5 in Supporting Information) is used to compute $m_{x}^{*}$ and $m_{y}^{*}$. The in-plane Young’s modulus, $C_{2\mathrm{D}}$, is defined as 55, 1:
@@
Recent advancements in understanding the intrinsic mobility of 2D TMDCs have emphasized the critical role of LO phonon scattering, in addition to the traditionally considered LA phonon scattering 25. To accurately capture these effects, many previous studies have utilized full electron-phonon coupling (EPC) calculations based on density functional perturbation theory (DFPT) combined with Wannier interpolation 56, 57, 58, 59. This approach enables precise computation of EPC matrix elements over dense k- and q-point grids, effectively overcoming the limitations of DPT, which assumes isotropic electron-phonon coupling and predominantly accounts for LA phonon scattering. However, while the EPC method offers a more accurate and comprehensive depiction of phonon-limited mobility, it is computationally intensive and demands substantial resources for high-precision simulations 60.
A more practical approach is to separately calculate the contributions of LO and LA phonon modes and subsequently determine the total mobility using Matthiessen’s rule 25. This can be achieved by employing the DPT formula for LA phonon scattering and the formula for Fröhlich interaction in polar 2D materials 27, 25. Including LO phonon scattering is particularly vital for polar materials, as its strength is closely linked to the Born effective charge, which quantifies the degree of electric polarization change caused by atomic vibrations. A larger Born effective charge results in stronger carrier scattering and, consequently, reduced mobility. This insight highlights the necessity of advancing beyond conventional DPT to accurately evaluate intrinsic mobility, especially for Janus-TMDC heterobilayers, where LO phonon scattering significantly influences carrier transport properties. The total scattering rate near the band edges can be written as
@@
$\begin{split}\frac{1}{\tau_{\text{LO}}}&=\frac{e^{4}\hbar n_{\hbar\omega}}{M_{M}M_{X1}M_{X2}S_{0}d_{\text{eff}}^{2}(\varepsilon\varepsilon_{0})^{2}(\hbar\omega)^{2}}\times\\ &~{}~{}~{}~{}~{}\left[\left|Z_{MB}\right|\sqrt{M_{X1}M_{X2}}+\left|Z_{X1}\right|\sqrt{M_{M}M_{X2}}+\left|Z_{X2}\right|\sqrt{M_{M}M_{X1}}\right]^{2},\end{split}$ (13)
@@
where $\tau_{\mathrm{LO}}$ represents the relaxation time, quantifies the rate at which charge carriers lose momentum due to phonon scattering, $e$ is the elementary charge, $n_{\hbar\omega}$ is the Bose-Einstein distribution at the phonon frequency at the gamma point, $A$ is the equilibrium area, $d_{\text{eff}}$ is the effective thickness, $\varepsilon$ is the in-plane optical dielectric constant, and $\varepsilon_{0}$ is the dielectric constant. Additionally, $M_{M}$, $M_{X1}$, and $M_{X2}$ are the atomic masses of the metal and the two chalcogens in the structure, while $Z_{MB}$, $Z_{X1}$, and $Z_{X2}$ are the Born effective charges of the metal and the two chalcogen atoms(see Table S2 in Supporting Information for reference). Thus, the mobility contributed by the LO phonon mode is given by
@@
In Janus-TMDC heterobilayers, two distinct LO phonon modes, LO1 and LO2, correspond to the vibrations of the TMDC and Janus layers, respectively, as shown in Figure 4. The calculated mobilities for $\mu_{\text{LO1}}$ and $\mu_{\text{LO2}}$ are provided in Table 3. Then the total mobility $\mu_{\text{total}}$ is then determined using Matthiessen’s rule as follows 25:
@@
The calculated $\mu_{\text{LO1}}$, $\mu_{\text{LO2}}$, and $\mu_{\text{total}}$ are listed in Table  3. Our results reveal that LO phonon scattering significantly limits mobility in Janus-TMDC heterobilayers, which leads to lower carrier mobility compared to those predicted by DPT alone (see Table  3). Importantly, our approach to calculating the mobility of TMDC materials and their Janus heterobilayers proves to be not only more precise but also computationally less expensive than methods relying solely on DPT or EPC. This balance between accuracy and efficiency underscores the robustness of our methodology in capturing the intricate phonon-electron interactions in these systems.
3.6 Enhanced mobility in Janus/TMDC heterobilayers
In the approach using the Fröhlich model to calculate the mobility of polar materials, Cheng and Liu 25 compared the mobilities of monolayer MoS2, MoSe2, WS2, and WSe2 calculated via Fröhlich scattering and the EPC method 25. Their findings revealed that the mobilities in these materials are primarily influenced by the LO and LA phonon modes. However, compared to the Fröhlich mobility of Janus/TMDC heterobilayers listed in Table 3, our heterobilayers exhibit significantly higher mobilities than those of monolayer TMDCs 25. This enhancement in mobility can be attributed to three main factors: (1) dielectric screening from the other layers, (2) remote phonon scattering from distant polar-optical phonons, and (3) changes in the Fröhlich scattering strength due to polar-optical phonon coupling 61.
Additionally, we find that in TMDC/Janus heterobilayers, the internal electric field modifies the Born effective charges, reducing their magnitude and thereby increasing the Fröhlich mobility. For instance, the Born effective charges of W atoms in monolayer WSe2 and WS2 are $-0.65$ and $-1.29$, respectively 25, whereas those in WS2-SMoSe, WSe2-TeWSe, and WSe2-SWSe heterobilayers are $-0.097$, $-0.52$, and $-0.57$, respectively. Another contributing factor is the effective thickness, $d_{\text{eff}}$. As shown in Eq. 13 and 14, the increasing $d_{\text{eff}}$ enhances the Fröhlich mobility. Consequently, the heterobilayers achieve higher mobilities compared to their monolayer counterparts. Furthermore, very recent corrections to the polar-optical phonon interaction in the Fröhlich model demonstrated that the Fröhlich mobility of InSe in an InSe/BN vdW heterostructure was enhanced by $2.5$ times compared to that of monolayers of InSe 61.
This significant mobility enhancement in heterobilayers compared to monolayers suggests that the Z-scheme alignment of our heterobilayers offers a new insight advantage: not only does it provide a favorable alignment for photocatalytic water splitting by reducing the electron-hole recombination rate, but it also significantly amplifies mobility, making them superior to monolayers.